OpenStudy (strawberry17):

Please show how to solve this: If √x - a = b, then x = ? (-a is not under the root sign)

OpenStudy (strawberry17):

Easy problem, but I just forgot how to solve it

zepdrix (zepdrix):

\[\large \sqrt x-a=b\] So let's start by adding \(\large a\) to each side. What do you get? :O

OpenStudy (strawberry17):

√x = a + b

zepdrix (zepdrix):

k looks good so far c:

OpenStudy (strawberry17):

does x = (a + b)^2

zepdrix (zepdrix):

Yes! Good! :) That might be your final answer, depending on whether or not your teacher wants you to expand out the binomial containing a and b.

OpenStudy (strawberry17):

okay, thanks so much for your help :)